Using the component method, solve for the resultant vector given the
following:
A = 250 km, 35^o E of N
B = 120 km, 60^o W of N

Answer by ikleyn(52874)

.
Using the component method, solve for the resultant vector given the
following:
A = 250 km, 35^o E of N
B = 120 km, 60^o W of N
~~~~~~~~~~~~~~~~~~~


            Make a sketch and mark the participating angles there.


We should present vectors A and B in component form, then add the components.


In component form, A = (Ax,Ay), where Ax is x-component of the vector A 
                                         (perpendicular to N-direction)

                                      Ay is y-component of the vector A
                                         (along N-direction)                              


    Ax = |A|*sin(35°) = 250*sin(35°),

    Ay = |A|*cos(35°) = 250*cos(35°).


Similarly for vector B: 

    Bx = -|B|*sin(60°) = -120*sin(60°),

    By = |B|*cos(60°) = 120*cos(60°).


Make all necessary calculations.


Then your resultant vector is the sum (A + B) 

    with x-component Ax + Bx perpendicular to N-direction and

    with y-component Ay + By parallel to N-direction.

You have all necessary instructions and explanations to complete the assignment on your own.